ADNAN MUHAMMED

Grade 12,

(1) Two particles, each of mass m and carrying charge Q, are separated by some distance.If they are in equilibrium under mutual gravitational and electrostatic forces,then Q/m (in C/Kg) is of the order of (a) 10 -5 (b) 10 -10 (c) 10 -15 (c) 10 -20

(1) Two particles, each of mass m and carrying charge Q, are separated by some distance.If they are in equilibrium under mutual gravitational and electrostatic forces,then Q/m (in C/Kg) is of the order of


(a) 10-5


(b) 10-10


(c) 10-15


(c) 10-20

Grade:11

5 Answers

Mujahid Ahmed
10 years ago

K Q2/r2 = Gm2/r2

Q2/m2 = G/K

= 6,6 x 10-11/9 x 109

= 10-20

 

Approve if i helped :)

Abhishekh kumar sharma
10 years ago

10^-15

anurag hundal
10 years ago

guess its c!

Jatin Singh
7 years ago
repulsive electrostatic force:                         attractive gravitational force:
F1=kQ^2/r^2                                                   F2=Gm^2/r^2
 
for equilibrium :F1=F2
kQ2/r2=Gm2/r2
=> Q/m = (6.6*10^-11/9*10^9)1/2
=> Q/m=0.8*10-10
Hence, ORDER OF Q/m=10-10
 
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
repulsive electrostatic force(since charges are same in magnitude and sign):            
F1 = kQ2/r2                                                 
attractive gravitational force:
F2 = Gm2/r2
for equilibrium :F1 = F2
kQ2/r2 = Gm2/r2
=> Q/m = (6.6*10-11/9*109)1/2
=> Q/m=0.8*10-10
Hence, ORDER OF Q/m=10-10
Hence (b) is the correct option.
 
Hope it helps.
Thanks and regards,
Kushagra